OpenStudy (anonymous):

Simplify: 7 radical 3 + 8 radical 3 minus 2 radical 2

OpenStudy (anonymous):

\[7\sqrt{3}+8\sqrt{3}-2\sqrt{2}\]

OpenStudy (anonymous):

you can add the radical 3's but not the 2's

OpenStudy (anonymous):

A. \[13\sqrt{5}\] B. \[13\sqrt{6}\] C. \[15\sqrt{3}-2\sqrt{2}\]

OpenStudy (anonymous):

so 11 rad 3, - 2rad 2

OpenStudy (anonymous):

sorry.. 7+8 = 15 :(

OpenStudy (anonymous):

so 15 rad 3 - 2 rad 2.... sorry about the # of posts

OpenStudy (anonymous):

Thanks! :) And it's no problem!! :)

OpenStudy (anonymous):

mmk. basically think of rad 3 as an "x" and rad 2 as a "y". you can add the x's but you cant mix with y's :)